Number 11 please. And please explain the final step to your y= equation

10–14 SERIES SOLUTIONS Find a power series solution in powers of x. Show the details. 10. y - y + xy = 0 11. y - y + x’y
Answer #1

. ). 9 у х* y = 0 #uu 9- 4ot ax +4 +ux +4° +а хаах +4, + + - .--.- у - 24,+ (0,1 + 12 * + 20 - 300 x + --- C-9 = -а, - 29 х- 9 - y +x o (242-4 ) + (саз-ге) х -бо, - зе, х* + (204, +94 + а) х* + (3sag tsag + 1 Equating the coeltidents on Its & RHSNow 20g - uanta, -o *200g = naural as= u( al ao sola 20 309-5a5 ta,=0 300g = sag az 3046 = 15 (saya) - al Therefore substitut+-- У= a +4 + 4 4 4 + + + 92 % +4х+аг х2 + а, « , -а) * 17an ач ( Ans